# SELFI: Autonomous Self-Improvement with Reinforcement Learning for Social Navigation

### v1

@@ V-B Performance analysis of the fine-tuned control policy
To evaluate how well our method behaves around humans, we also conduct additional experiments with ten human subjects. We recruited ten subjects from among graduate students, visiting scholars, and staffs on campus. We consider the balance to be as diverse in gender (5 male, 5 female), professional (6 student, 4 non-student), and origin (4 North America, 4 Asia, 1 Europe, 1 Latin America) as possible. We asked the subjects to walk through the space during a robotic navigation trial. Subjects were not told which method was being used for each trial. After each experiment, subjects were asked qualitative questions with a 5-point Likert scale: (1) was the robot motion intrusive? (2) was the robot friendly and natural? (3) did the robot motion feel smooth? (4) did the robot pre-emptively move out of your way?
Figure 7: Evaluation of socialness by human rating. The height of each bar indicates the mean and the range line indicates the standard deviation. Larger is better in all ratings. $*{}^{*}$ indicates the statistical significance of our method on the t-test with $p<$ 0.05.
TABLE I: Closed-loop Evaluation of trained control policies. We evaluate our proposed method with some baselines. CP is the number of collision for pedestrians, CO is the number of collision for the tiny objects, Int is the number of interventions by teleoperators, PSV is personal space violation duration, NCO is near-collision duration, and UFS is duration on uneven floor surface.
Method SPL $\uparrow$ STL $\uparrow$ CP $\downarrow$ [#] CO $\downarrow$ [#] Int $\downarrow$ [#] PSV [s] $\downarrow$ NCO [s] $\downarrow$ UFS [s] $\downarrow$
Residual RL 0.817 0.641 3.200 4.333 2.067 28.731 12.365 8.073
SACSoN with fine-tuning 0.838 0.596 0.800 2.667 0.733 18.018 11.229 3.578
Ours 0.918 0.739 0.200 1.400 0.267 7.978 3.938 3.067
Figure 8: Visualization of the robot behavior when interacting with the pedestrian. Robots and pedestrians with the same color indicate the same time. The black dotted lines indicate the robot and pedestrian at the time of closest proximity.
Figure 9: Visualization of the robot behavior of collision avoidance for small obstacles and uneven floor mat.
Figure 10: The number of intervention on online learning in three different environments. The lines indicate the mean and the areas indicate the standard deviation.
Figure 7 shows the means and standard deviations of the scores from this study. Note that we flip the score for the first question such that higher is better for the entire bar plot. Residual RL often penetrates the personal space and occasionally collides with the human subject. As the result, Residual RL has the worst score. SACSoN with fine-tuning shows similar behavior in navigation and the closest score to our method. However, our method performs better overall. The statistical significance of our method is confirmed in three questions, (1), (2) and (4) on the t-test with $p<$ 0.05.
Figure 8 visualizes the robot behavior when the pedestrian and robot pass each other. In these time lapse illustrations, the color of the pedestrian and robot indicates the timestep (i.e., a yellow pedestrian and a yellow robot indicate the same point in time). In addition, we use a black dotted line for the robot and the pedestrians when robots and people are in closest proximity. In both cases, the robot with the baseline method (right) gets stuck in front of the pedestrians and penetrates the personal space for an extended length of time. Especially in case A, the baseline fails to reach the goal position and collides with the chair with thin legs. With our method, an evasive maneuver is initiated at an early stage and succeeds in smoothly passing a pedestrian without getting stuck (left). Although the robot is close to the pedestrians when passing, our control policy minimizes how long the robot penetrates the pedestrian’s personal space.
Figure 9 shows the time lapsed images when avoiding the small unseen objects and the uneven rubber mat. Our method naturally avoids collision with the small objects, though this presents a challenge for the initial SACSoN policy. In addition, our methods avoids traveling on the uneven mat by giving negative reward in $C_{s}$ on online training. Please see our supplemental materials for more details.
V-C Interventions during online training
Reducing human interventions is important in online training to enable autonomous adaptation in the real world. However, it is known that the data distribution shift causes the performance degradation leading navigation failures on fine-tuning and leads a lot of interventions to keep online learning in navigation. To evaluate the online training process, we count the number of human interventions in each navigation loop for Q2.
V-D Pre-training with data from the target environment
When robots operate continuously in the same environment, they can directly collect and utilize a large amount of in-domain data. In the last evaluation, we show that a large offline dataset of interactions from the target environment can boost the performance of our method during online training for Q3. Specifically, we use offline RL to train our actor and critic on the collected dataset from the target environment on top of SACSoN. Then, we fine-tune both networks on online training by SELFI.
During offline RL training, we leverage the collected dataset from our study. While finding good hyperparameter for online learning, we collected a dataset of 50 hours of experience in Environment 1. Figure 11 shows the number of intervention for each loop in Environment 1. Even at beginning, our method leverages a large dataset, Ours + prior data only needs one intervention and can navigate the robot without any interventions after 5 laps, because of faster training with the pre-trained actor and critic on the large dataset. Similar to Table I, Table II shows the mean value of the selected metrics on 5 laps in the Environment 1. Ours + prior data shows a remarkable gap against Ours in every metric except SPL. Since Ours + prior data takes a larger deviation from the original path to avoid violating the personal space of close pedestrians (to decrease PSV), SPL is slightly worsen.
@@ V-D Pre-training with data from the target environment
Figure 11: The number of intervention on online learning in Environment 1. Ours + prior data indicates our method leveraging large dataset in same environment. Residual ${}^{\dagger}$ uses the SACSoN policy as $\pi_{\text{base}}$.
TABLE II: Closed-loop Evaluation of trained control policies at Environment 1. Ours + prior data indicates our method leveraging large dataset in same environment. Residual ${}^{\dagger}$ uses the SACSoN policy as $\pi_{\text{base}}$.
Method SPL $\uparrow$ STL $\uparrow$ CP $\downarrow$ [#] Int $\downarrow$ [#] PSV [s] $\downarrow$
[…]
FastRLAP 0.702 0.306 2.000 10.000 43.623
Residual RL 0.780 0.568 4.20 3.000 21.754
Residual ${}^{\dagger}$ 0.598 0.466 2.800 9.400 39.627
SACSoN with fine-tuning 0.788 0.549 0.200 1.000 19.534
Ours 0.932 0.699 0 0.600 10.712
Ours + prior data 0.900 0.713 0 0.200 6.610
VI Discussion
In this paper, we proposed an online self-improving method, SELFI, to quickly fine-tune a control policy pre-trained with model-based learning. SELFI combines model-based learning and model-free RL its training objectives to take advantage of the best parts of both approaches. The same objectives used in offline learning are introduced into online learning to stabilize the learning process. The performance of the pre-trained control policies are improved via Q-functions from online model-free RL.
In the evaluation, SELFI was implemented to fine-tune the SACSoN policy [18] for vision-based navigation. SELFI enables us to quickly learn complex robotic behavior, such as pre-emptive collision avoidance for pedestrians, collision avoidance for the small and transparent obstacles, and preferences for traversing on smooth surfaces. These behaviors are difficult to learn on offline training due to the modeling errors and data distribution shift. In addition, compared to various baseline methods, SELFI did not require much human intervention during online learning. The performance of the trained control policy by SELFI is also visualized in our project page.
While our method enables us to quickly fine-tune a pre-trained control policy, it has some limitations. For effective online learning, the balance between objectives from model-based learning and the learned Q-function is important, but this balance cannot be predicted in advance and requires some trial-and-error with real robots. Although the reward for socialness in navigation is given only for personal space violations, human-in-the-loop online learning with human evaluations can lead to better behavior.
Acknowledgments
This research was supported by Berkeley DeepDrive at the University of California, Berkeley, and Toyota Motor North America. And, this work was partially supported by ARL DCIST CRA W911NF-17-2-0181 and ARO W911NF-21-1-0097. We thank Qiyang Li and Mitsuhiko Nakamoto for advising the mathematical formulation. We would like to express their gratitude to Roxana Infante, Ami Katagiri, Pranav Atreya, Stephan Allenspach, Lydia Ignatova, Toru Lin, Charles Xu, Kosuke Tahara, Katsuhiro Kutsuki and Catherine Glossop for their valuable assistance in evaluating SELFI.
Appendix
VI-A Robot pose estimation with AR markers
@@
For localization, we mount a tracking camera on our robot that measures visual odometry. However, the shaky robot motion and insufficient visual features deteriorate the accuracy of visual odometry. To get a more accurate estimate of the robot’s global position, we place one AR marker every 15 [m] along the robot’s trajectory and suppress the localization error for stable online learning.
@@
Hence, we can obtain $T^{c}_{i}$ by calculating $T_{i}^{g}\cdot T_{i}^{AR_{g}}\cdot(T^{vo}_{t}\cdot T_{i}^{AR_{r}})^{-1}$. During online learning, we update $T^{c}_{i}$ every time we detect an AR marker, and we use $T^{c}_{i}$ to calculate $T$ until the robot detects the next AR marker. The robot pose $p$ is uniquely calculated from the robot pose matrix $T$. And, the goal pose $\hat{g}$ on the robot local coordinate can be uniquely calculated from $T^{-1}T_{g}$, where $T_{g}$ is the global goal pose matrix for the goal image $I^{g}$.
VI-B Details of network structure

### v2

@@ Appendix E Pre-training with data from the target environment
Figure 9: The number of intervention on online learning in Environment 1. Ours + prior data indicates our method leveraging large dataset in same environment. Residual RL uses the SACSoN policy as $\pi_{\text{base}}$. The horizontal axis indicates the number of times the robot laps the loop.
Table 2: Closed-loop Evaluation of trained control policies at Environment 1. Ours + prior data indicates our method leveraging large dataset in same environment. Residual RL uses the SACSoN policy as $\pi_{\text{base}}$. * indicates the use of the ground truth goal pose to generate velocity commands.
Method IDV [s] $\downarrow$ CP $\downarrow$ [#] Int $\downarrow$ [#] SPL $\uparrow$ STL $\uparrow$
Sampling-based motion planning* 23.909 1.800 0.400 0.839 0.678
@@
FastRLAP 43.623 2.000 10.000 0.702 0.306
Residual RL 21.754 4.200 3.000 0.780 0.568
Residual RL 39.627 2.800 9.400 0.598 0.466
SACSoN with fine-tuning 19.534 0.200 1.000 0.788 0.549
Ours 10.712 0 0.600 0.932 0.699
Ours + prior data 6.610 0 0.200 0.900 0.713
Appendix F Robot pose estimation with AR markers
@@
For localization, we mount a tracking camera, Intel T265 on our robot that measures visual odometry. However, the Intel T265 can not maintain the same globalframe before and after rebooting due to battery replacement. In addition, the shaky robot motion and insufficient visual features deteriorate the accuracy of visual odometry. To have same global frame and to get a more accurate estimate of the robot’s global position, we place one AR marker every 15 [m] along the robot’s trajectory and suppress the localization error for stable online learning.
@@
Here $T$ is on the same global coordinate as $T_{i}^{g}$. Hence, we can obtain $T^{c}_{i}$ by calculating $T_{i}^{g}\cdot T_{i}^{AR_{g}}\cdot(T^{vo}_{t}\cdot T_{i}^{AR_{r}})^{-1}$. During online learning, we update $T^{c}_{i}$ every time we detect an AR marker, and we use $T^{c}_{i}$ to calculate $T$ until the robot detects the next AR marker. The robot pose $p$ is uniquely calculated from the robot pose matrix $T$. And, the goal pose $\hat{g}$ on the robot local coordinate can be uniquely calculated from $T^{-1}T_{g}$, where $T_{g}$ is the global goal pose matrix for the goal image $I^{g}$.
While we use AR markers due to our design decision using only an RGB camera and limitations of the Intel T265, it is important to note that the usage of AR markers is not mandatory. Our system setup serves as just one example. If potential users relax the restriction of using only the camera, they can leverage other sensors such as GPS, LiDAR, and/or depth cameras to estimate the robot’s pose without requiring AR markers. It’s worth mentioning that we can run our system without AR markers during inference using other vision-based localization techniques [61, 21, 74, 77, 25], but we use AR markers to identify the corresponding goal image ID, focusing on evaluating the trained control policy.
Appendix G Details of sampling-based motion planning
We implemented sampling-based motion planning as the baseline to bridge the learning-based approach with broader robotic motion planning. We generated 15 motion primitives assuming steady linear and angular velocity commands for 8 steps (2.664 s), which is the same horizon as our method and the strongest baseline, SACSoN. The pairs of linear and angular velocity commands are $(v_{s},\omega_{s})=$ (0.0, 0.0), (0.2, 0.0), (0.2, 0.3), (0.2, 0.6), (0.2, 0.9), (0.2, --0.3), (0.2, --0.6), (0.2, --0.9), (0.5, 0.0), (0.5, 0.3), (0.5, 0.6), (0.5, 0.9), (0.5, --0.3), (0.5, --0.6), (0.5, --0.9). We selected these 15 motion primitives by balancing computational load and navigation performance.
[…]
Here, $p^{g}_{i_{c}+1}$ indicates the next subgoal pose. The first term on the right-hand side calculates the squared errors between all 8 poses in the $j$-th motion primitive and the goal pose and selects the minimum one to evaluate the goal-reaching performance. $C_{ob}$ is a constant value used to filter out trajectories that collide with static obstacles. We calculate $C_{ob}$ as follows:
$C_{ob}$ $=$ $\begin{cases}1000.0\hskip 14.22636pt(\mbox{if}\hskip 4.2679ptd_{s}<r_{r})\\ \hskip 8.53581pt0.0\hskip 14.22636pt(\mbox{otherwise})\\ \end{cases},$ (7)
where $d_{s}$ is the minimum distance between all 8 poses in the $j$-th motion primitive and the estimated point clouds corresponding to the static obstacles. Similar to SACSoN [5] and ExAug [25], a collision is determined when the distance between the static obstacle and the robot is less than the robot’s radius $r_{r}$. To ensure fair evaluation with other methods in vision-based navigation that only uses an RGB camera, we utilize estimated point clouds from the current observation of the RGB camera. Additionally, $C_{ped}$ is a constant value used to filter out trajectories that violate the intimate distance with pedestrians. Following SACSoN [5], we predict the future trajectory of pedestrians as $\{p^{ped}_{i}\}_{i=1\ldots 8}$ and assess whether each motion primitive violates the intimate distance or not.
$C_{ped}$ $=$ $\begin{cases}1000.0\hskip 14.22636pt(\mbox{if}\hskip 4.2679ptd_{ped}<0.5+r_{r})\\ \hskip 8.53581pt0.0\hskip 14.22636pt(\mbox{otherwise})\\ \end{cases}.$ (8)
Here, $d_{ped}=\min_{i}\text{dist}(p^{ped}_{i},^{s}p_{i}^{j})$, and $\text{dist}()$ is the function used to calculate the distance between two poses. The values of $C_{ob}$ and $C_{ped}$ are set to 1000.0, which is much larger than the first term on the right-hand side in Eqn 6, to filter out inappropriate motion primitives during the selection process. We choose the motion primitive with the minimum $J_{s}^{j}$ and assign the corresponding velocity commands $v_{s}$ and $\omega_{s}$ to control the robot during navigation. Note that we utilize the same point cloud estimator as well as the pedestrians’ trajectory predictor as our method and SACSoN to have fair comparison.
Figure 11 shows the implemented sampling-based motion planning. Since pedestrians are not present in [b], this baseline selects the motion primitive that moves toward the goal position. However, in [a], where the pedestrian’s predicted trajectory crosses between the robot’s current position the goal position, the baseline selects a motion primitive that maintains a sufficient distance from the pedestrian’s predicted trajectory. After the pedestrian passes through the area, the baseline will select a trajectory towards the goal. In Fig. 11, the blue lines with the “x” markers are the motion primitives and the black dots are the estimated point clouds.
Figure 11: Overview of sampling-based motion planning.
Appendix H Details of network structure